OpenStudy (anonymous):

turn 10sqrt(10)+11sqrt(11) into the following exact answer: sqrt( 2331+ 220sqrt(110))

OpenStudy (anonymous):

\[10\sqrt{10}+11\sqrt{11} = \sqrt{2331+220\sqrt{110}}\]

OpenStudy (anonymous):

\( 10\sqrt{10}+11\sqrt{11} = \sqrt{ (10+11)^2 +11(10+10)^2 \sqrt{(11\times 10)}}\)

OpenStudy (anonymous):

\((10\sqrt{10}+11\sqrt{11})^2 = (10\sqrt{10})^2+(11\sqrt{11})^2+2\times 10\sqrt{10}\times 11\sqrt{11} \\= 10^2.10+11^2.11+ 2.10.11.\sqrt{11\times 10}\\ =(10^3+11^3)+(2\times10\times11)\sqrt{11\times 10} \\ =2331+220\sqrt{110 }\)

OpenStudy (anonymous):

hmmm note that 220=11*(10+10) not 11*(10+10)^2

OpenStudy (anonymous):

Your answer should be under a square root. Your answer = 3026.70... while the problem equals 68.105....

OpenStudy (anonymous):

no no please read my answer \((10\sqrt{10}+11\sqrt{11})^2 = 2331+220\sqrt{110 } \) now take root for both sides \(\sqrt{(10\sqrt{10}+11\sqrt{11})^2} = \sqrt{2331+220\sqrt{110 } }\\ 10\sqrt{10}+11\sqrt{11} = \sqrt{2331+220\sqrt{110 } } \)